sakuragi is spinning a basketball with a radius of 0.15 m on the tip of his finger . Determine the mass of the ball if it's moment of inertia is five point 568 x 10 negative 3 kg x m squared

1
Expert's answer
2021-12-01T10:29:02-0500

Let's model the ball with a hollow sphere. Then its moment of inertia is given as follows:


"I = \dfrac23mr^2"

where "m" is the mass of the ball and "r = 0.15m" is its radius. Since "I = 5.568\times 10^{-3}kg\cdot m^2", find:


"m = \dfrac{3I}{2r^2}\\
m = \dfrac{3\cdot 5.568\times 10^{-3}}{2\cdot 0.15^2} \approx 0.37kg"

Answer. 0.37 kg.